OpenStudy (anonymous):

1/18x^2-1/9x=1 Use the quadratic formula to solve.

hartnn (hartnn):

hello :) have you tried this ?

OpenStudy (anonymous):

I have tried. I got \[(\sqrt{19}+1)(-\sqrt{19}+1)\]

hartnn (hartnn):

did you miss 1/18 ?? because i get, x= (1/18) (sqrt 19-1) and x=(1/18)(-sqrt 19-1)

OpenStudy (anonymous):

That might be what I'm doing wrong

hartnn (hartnn):

ask if any more doubts...

OpenStudy (anonymous):

Thanks (: